i would like to know how to calculate the capacitance due to two concentric hollow spherical shells of radii r1 and r2 with r1each having
1.equal charge
2.unequal charge

Grade:12

1 Answers

Askiitians Expert Soumyajit IIT-Kharagpur
14 years ago

Dear STUDENT NAME,

Ans:- The thing is that if you want to calculate the capacitance then first of all you have to give it some charge. Now let's give a charge Q to it. and it is given to the inner sphere. As from the Gauss's law we know that there can not be any charge inside a hollow sphere and hence the entire charge will come out to the external surface of the outer sphere.So the potential of the inner sphere will be V1=Q/Kr1 -Q/kr2+Q/kr2 = Q/Kr1..............(1)                                                           

(Because the charge appearing in the inner and outer surface of the outer sphere is Q && -Q respectively. Thus to calculate the potential in the inner sphere, we must consider all the charges.

But if we consider the outer hollow sphere then the total charge inside this is 0 hence it's potential is V2=Q/Kr2 .....(2)

potential difference = V1-V2=Q/K(1/r1 - 1/r2)=ΔV

Now the capacitance is Q/ΔV=Kr1r2/(r1 - r2)

Hence C=Kr1r2/(r1-r2)

Please feel free to post as many doubts on our discussion forum as you can. If you find any question
Difficult to understand - post it here and we will get you the answer and detailed solution very quickly. We
are all IITians and here to help you in your IIT JEE preparation.


All the best STUDENT NAME !!!

 


Regards,

Askiitians Experts
Soumyajit Das IIT Kharagpur